OpenStudy (mckenzieandjesus):

Please help me! (: write an equation in slope-intercept form (y=mx+b) of the line through point P(6,7) with slope -2 A. y-7=-2(x-6) B. y=-2x+7 C. y=-2x+19 D. y-6=-2(x-7)

OpenStudy (mckenzieandjesus):

@LegendarySadist

pooja195 (pooja195):

\[\huge~y-y1=m(x-x1)\] m=slope (x1,y1)

OpenStudy (mckenzieandjesus):

so 6-61=-2 (7-71) did i do it right?

pooja195 (pooja195):

why did u keep the ones there?

OpenStudy (mckenzieandjesus):

oh oops. so the answer would be y-6=-2(x-7)?

pooja195 (pooja195):

good now can you solve that?

OpenStudy (mckenzieandjesus):

Well isn't that all I had to do the answer would be D right?

pooja195 (pooja195):

\(\color{blue}{\text{Originally Posted by}}\) @mckenzieandjesus Please help me! (: write an equation in slope-intercept form (y=mx+b) of the line through point P(6,7) with slope -2 A. y-7=-2(x-6) B. y=-2x+7 C. y=-2x+19 D. y-6=-2(x-7) \(\color{blue}{\text{End of Quote}}\) Read the directions again :)

pooja195 (pooja195):

do u see it ?

OpenStudy (mckenzieandjesus):

Well its multiple choice and those were the answers I was given to choose. I thought I just need to put (6,7) and the slope -2 into a equation..

pooja195 (pooja195):

It wants it in slope intercept form :) y=mx+b \[\huge~ y-6=-2(x-7)\] Distribute \[\huge~-2(x-7) \]

OpenStudy (mckenzieandjesus):

Ohhh... so the answer would be B! oops lol. thank you so much!

pooja195 (pooja195):

yep :)

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (mckenzieandjesus):

can you help me with another question?